### v3

# Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications

Abstract
Tiny machine learning (TinyML) for low-power devices lacks robust datasets for development. We present Wake Vision, a large-scale dataset for person detection that contains over 6 million quality-filtered images. We provide two variants: Wake Vision (Large) and Wake Vision (Quality), leveraging the large variant for pretraining and knowledge distillation, while the higher-quality labels drive final model performance. The manually labeled validation and test sets reduce error rates from 7.8% to 2.2% compared to previous standards. In addition, we introduce five detailed benchmark sets to evaluate model performance in real-world scenarios, including varying lighting, camera distances, and demographic characteristics. Training with Wake Vision improves accuracy by 1.93% over existing datasets, demonstrating the importance of dataset quality for low-capacity models and dataset size for high-capacity models. The dataset, benchmarks, code, and models are available under the CC-BY 4.0 license, maintained by the Edge AI Foundation.
1 Introduction
Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 38, 1, 6]. Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints. For instance, the models must often be within hundreds of kilobytes of memory in size, nearly four orders of magnitude less than mobile phones[2].
TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [5, 37] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models. This creates a fundamental challenge where we can not develop robust TinyML models due to existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
Table 1: A comparison of person detection and image classification datasets.
Total # of Person Images Fine-Grained Suitable for
Dataset Images Train Validation Test Filtering TinyML
Our work – Wake Vision (Quality) 1,322,574 624,115 9,291 27,881
Our work – Wake Vision (Large) 5,760,428 2,880,214 -- --
Visual Wake Words [5] $123,287$ $36,000$ $3,926$ $19,107$
CIFAR-100 [20] $60,000$ $2,500$- $500$
PASCAL VOC 2012 [8] $11,530$ $1,994$ $2,093$-
To address this critical gap, we present Wake Vision, a large-scale dataset specifically designed for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task that, while simple, enables crucial applications ranging from occupancy detection [31] and smart HVAC/lighting [39] to acting as an always-on ’wake model’ in larger ML systems [15].
Wake Vision advances the state-of-the-art by providing 6M images, approximately 100x more than the current leading dataset Visual Wake Words (VWW) [5], while achieving a 1.93% accuracy improvement. The dataset is derived from Open Images v7 [21, 19] and is released under a permissive CC-BY 4.0 license to facilitate broad adoption.
A key contribution of our work is the first systematic investigation of how dataset quality versus quantity impacts TinyML model performance. This question is particularly relevant for resource-constrained models, as prior work suggests that large, overparameterized models can effectively identify and adapt to training set errors [40, 9]. TinyML models, with their limited capacity, may exhibit fundamentally different behavior. To explore this hypothesis, we introduce two complementary training sets within the Wake Vision visual corpus: Wake Vision (Large), emphasizing scale, and Wake Vision (Quality), prioritizing label accuracy. Our experiments reveal that while Wake Vision (Quality) outperforms its larger counterpart in naive training scenarios, optimal performance is achieved through a training strategy that leverages both sets — using Wake Vision (Large) for pre-training followed by using Wake Vision (Quality) fine-tuning. To further explore the dataset quality and quantity trade-off for tinyML, we train a range of ResNet [14] style models across Wake Vision datasets of varying error rates.
Our experiments collectively reveal that small models are more sensitive to higher error rates than large models and that large models can benefit from larger datasets despite a higher error rate. These findings provide insight into the unique challenges of training highly constrained models.
Beyond raw performance metrics, real-world deployment of TinyML systems often demands robustness across diverse and challenging use case conditions. Internet-sourced datasets often underrepresent challenging scenarios, such as low-lighting conditions, leading to overly optimistic performance estimates based on standard test set metrics.
To enable a rigorous evaluation of real-world capabilities, we additionally develop a comprehensive suite of five fine-grained benchmarks based on our Wake Vision dataset. Our benchmarks systematically evaluate performance across multiple dimensions: distance (detection at varying ranges), lighting (performance in well-lit and poorly-lit environments), depictions (handling of drawings and sculptures), perceived gender (investigating potential gender biases), and perceived age (examining effectiveness across age groups).
This benchmarking framework addresses a gap in current TinyML assessment methods and enables researchers to identify potential biases, limitations, and areas for improvement that might be masked by aggregate metrics alone. We find that Wake Vision models outperform VWW on 13 of the 16 fine-grained sets, demonstrating robust performance further illustrating the advantages of a large diverse training dataset.
In summary, we present the following contributions:
  • A production grade computer vision dataset for TinyML containing over 6M images (100$\times$ larger than state-of-the-art) for person detection, with comprehensive filtering, permissive licensing, and accessibility through standard ML frameworks.
  • A cleaned test set with 2.2% label error rate, enabling more reliable evaluation of TinyML models.
  • A benchmark suite for systematically evaluating model robustness across five critical dimensions (perceived gender, age, distance, lighting, depictions), exposing potential failure modes in real-world deployments.
  • Evidence that TinyML models are more sensitive to dataset quality than larger models, providing guidance for dataset curation in resource-constrained settings.
  • A complete, maintained dataset infrastructure including benchmarks, evaluation code, and reproducible filtering pipeline to support ongoing TinyML research.
(a) Older Person
(b) Near Person
(c) Bright Image
(d) Female Person
(e) Depicted Person
Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
2 Background and Related Work
Person Detection and Its Role in TinyML
Person detection is a fundamental computer vision task whose job is to determine the presence or absence of humans in visual data. Unlike more complex tasks such as pose estimation or instance segmentation, person detection focuses on the binary classification problem.
This seemingly simple task has emerged as a canonical use case for TinyML systems [3, 1], driven by both theoretical foundations and practical considerations. In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility. Person detection represents an optimal compromise. It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications.
The significance of person detection extends beyond its computational characteristics, serving as a critical selective activation mechanism in resource-constrained systems. These lightweight models efficiently manage the activation of power-intensive sensors and processors by determining human presence. This architectural pattern has demonstrated particular value in diverse applications, including occupancy detection [31], smart HVAC systems [39], and as an always-on "wake model" in larger ML pipelines [15]. Moreover, this approach enables substantial energy savings while preserving privacy through on-device processing.
Existing Person Detection Datasets
In the TinyML computer vision domain, the VWW dataset [5, 6, 24, 2, 1] has established itself as the de facto standard. It remains the only open-source dataset specifically designed for person detection with direct commercial applicability and potential for real-world impact. However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [25]) restrict its accessibility and utility.
Although other data sets contain person labels within general image classification collections, such as Cifar-100 [20] and PASCAL Visual Object Classes [8], these present significant drawbacks for TinyML applications. Specifically, their inadequate representation of the open “no-person” class can lead to poor perceived performance of TinyML models [5]. Our Wake Vision dataset addresses these limitations by providing nearly two orders of magnitude more images than any existing public, permissively licensed dataset (Table 1). Furthermore, it distinguishes itself as the only person detection dataset offering a fine-grained benchmark suite and official distribution through popular services like TensorFlow Datasets (TFDS) and Hugging Face Datasets.
3 Wake Vision: Dataset and Benchmark Suite
We introduce Wake Vision, a new dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is almost two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. Fig. 1 illustrates examples of images in the dataset. In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the data set and comparing it with VWW Dataset.
3.1 Label Generation
A large person detection dataset is indispensable for TinyML research. It can, however, be challenging to obtain. Manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images [21, 19], which contain both image-level and bounding box labels of various objects. Image-level labels describe which objects are present in an image but not localized to a specific point in the image. Bounding box labels provide a set of four coordinates forming a box around the object.
The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
Users can adapt Wake Vision to meet the needs of specific use cases by adding several configurable options to the open-source dataset creation script. For example, a configuration option exists to change whether depictions are labeled as non-persons or excluded from the dataset. We refer to Appendix I for details about these configuration options.
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1). The Open Images validation and test sets are fully labeled with bounding box labels. Therefore, both the Wake Vision validation and test sets are created using the higher-quality bounding box labels.
Large Training Set from Image-Level Labels
Image-level labels in Open Images contain little additional information than the raw labels. The only additional relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0 to 1 for labels generated by machines and is either 0 or 1 for labels negatively or positively verified by humans, respectively. See Sec. 3.2 for more information. We use this property to exclude images from Wake Vision if they contain only low-confidence person labels.
Quality Training Set From Bounding Box Labels
Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans. This ensures that bounding box labels are less likely to be false positives. Bounding box labels can be used to calculate an approximation of the area taken up by a person. We use this as a proxy for the distance of a person in the image, and by default exclude images where all person labels are far away, i.e., take up less than 5% of the image. An example of an image with a faraway person can be found in Fig. 2(a). Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction. See an example of an image containing a person with the depiction flag set in Fig. 2(b). By default, we use this flag to make sure that we do not consider a depiction a person. See Appendix C for an illustration of the bounding box filtering process.
(a) An example of a person far away
(b) An example of a depicted person
Figure 2: Examples of challenging outlier images.
3.2 Label Correction
Label errors are a challenge in computer vision that limit progress and obscure true performance [30, 4]. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision. All labels in the Open Images dataset originate from machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking. Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset can be used to accurately evaluate a model’s performance.
We use the Scale Rapid tool from Scale AI (https://scale.com/) to relabel the Wake Vision validation and test sets. Details about the relabelling process, including instructions given and total cost, can be found in Appendix N.
With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [29]. We perform an initial exploration of these methods in appendix O.
We report our estimated label error rate after label corrections versus VWW in Table 2. These estimates are all based on a random subset of 500 samples from each dataset, which the authors manually relabelled according to whether we saw a non-depiction person or body part in the image. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 2: Label error rate of VWW and Wake Vision after label correction. The Wake Vision error rate is broken down by set. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
Dataset Wake Vision Set Label Error
Visual Wake Words Train, Val, & Test 7.8%
Wake Vision Train (Large) 15.2%
Wake Vision Train (Quality) 6.8%
Wake Vision Val & Test 2.2%
3.3 Fine-grained Benchmark Suite
While Wake Vision provides a substantial improvement in scale and quality over existing datasets, we recognize that overall test set performance alone may not capture the nuanced challenges of real-world deployment for TinyML applications. Internet-sourced images typically exhibit a distinct bias towards well-composed photographs—they are often well-lit, properly framed, and feature clearly visible subjects. This stands in stark contrast to the challenging conditions where TinyML person detection systems are typically deployed, such as varying lighting conditions, extreme viewing distances, or partial occlusions [32].
To bridge this gap between benchmark performance and practical utility, we augment Wake Vision with a comprehensive fine-grained benchmark suite. This suite evaluates model robustness across challenging real-world scenarios where traditional accuracy metrics might mask significant failure modes. The benchmarks assess performance across critical dimensions, including lighting conditions, subject distance, and demographic attributes, enabling developers to identify potential biases or limitations during the design phase rather than after deployment. By providing these targeted evaluation sets, we enable a more nuanced analysis of model behavior beyond aggregate accuracy metrics.
The suite consists of five fine-grained benchmark sets. Each of the five fine-grained benchmarks has been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit compared to the inverse setting. Appendix D contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
Perceived Gender and Age
Table 3: Accuracy on the Wake Vision (WV) and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
Train
VWW Wake Vision (Quality) WV (Combined)
Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
WV 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [16]. This benchmark, generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [34], evaluates a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on perceived gender and age representation and are not necessarily representative.
Distance
This tests how the distance of people in images impacts model performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the faraway dataset will be more informative than its performance on the top-level test set. We created three data sets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
Lighting
The lighting data sets test the performance of Machine Learning (ML) systems in different lighting conditions. In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine-grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, a simple but effective method for distinguishing lighting conditions [41]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
Depictions
A particularly challenging task for a person detection model is to correctly reject depictions of people. In many use cases, a person detection model should not trigger on a depiction. For example, a room occupancy detector could incorrectly identify a painting on the wall as a person. This benchmark measures a model’s accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images that do not contain a depiction of any kind. Depictions of people can range from photo-realistic to crude stick figures.
3.4 Usability
Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [36] and HuggingFace Datasets [22] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
4 Results
In this section, we present a comprehensive evaluation of Wake Vision through five complementary analyses. First, we provide the first insights into best practices for Wake Vision by comparing its two training sets, which also provide initial insights into the trade-off between dataset quality and size in a TinyML setting. Second, we show that Wake Vision serves as an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models. Third, we validate Wake Vision’s benefits across diverse model architectures commonly used in the TinyML community, from MCUNet to ColabNAS families. Fourth, we thoroughly investigate the critical trade-off between dataset quality and size for TinyML models, providing insights into how model capacity influences sensitivity to label noise versus dataset scale. Finally, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios. Throughout these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
4.1 Training Set Evaluation
Table 4: Accuracy on test set of a MobileNetV2-0.25 trained on image-level and bounding box labels, with and without distillation. A MobileNetV2-1.0 model trained on the bounding box set is used as the teacher in KD.
Training set Label Source Training Set Size Accuracy Distilled Accuracy
Wake Vision (Large) Image-Level 5,760,428 80.80$\pm$0.1885.00$\pm$0.20%
Wake Vision (Quality) Bounding Box 1,248,230 84.89$\pm$0.1185.17$\pm$0.18%
Wake Vision (Combined) Both 5,760,428 85.72$\pm$0.04-
To evaluate the performance of the two Wake Vision training sets, we train identical models on each set. We use MobileNetV2 [33] models with a width modifier of 0.25 for 200,000 steps on 224x224x3 images using AdamW [26]. We use a lr of 0.002 with a cosine decay and a weight decay of 4E-6. After training, we compare each model’s performance on the common test set.
The result of this comparison is shown in Tab. 4. These results show that the Wake Vision (Quality) training set outperforms the Wake Vision (Large) training set by 4.09% test accuracy. This indicates that label quality is more important than quantity in this setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
Using the two training sets in unison to train a model where Wake Vision (Large) acts as a pre-training set, and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%. The two training sets together provide a foundation for further research on data-centric TinyML [27, 28].
4.2 Wake Vision & VWW Cross Evaluation
A key challenge in advancing TinyML research is ensuring that improvements can be readily adopted by the community without requiring significant modifications to existing workflows or architectures. While novel datasets can offer better quality or scale, their practical impact is limited if they require substantial changes to model architectures, training pipelines, or deployment processes. Therefore, we designed Wake Vision to serve as a drop-in replacement for VWW.
To validate this drop-in capability and demonstrate Wake Vision’s effectiveness, we do a cross-evaluation study. We train two identical MobileNetv2 [33] models according to the same training recipe as in Sec. 4.1. One model uses VWW’s training set, and the other uses Wake Vision’s Quality training set. After training these identical models using the respective datasets’ training sets with the same recipe for an equal number of steps, we evaluate their performance on the corresponding test sets. Creftab:wv-vww-cross-val shows the result of the cross-evaluation. The Wake Vision trained model outperforms the VWW trained model with a 0.26% improvement on VWW’s own test set, which indicates that Wake Vision is a direct improvement over VWW, not simply a domain shift.
Also, we achieve a 1.1% improvement over VWW on Wake Vision’s Test set, indicating that the new test set is more challenging. Table 3 also shows the performance of the Wake Vision (Combined) model described in Sec. 3.1. This model achieves 0.75% improvement over the Wake Vision (Quality) trained model on the VWW test set and 0.83% improvement on the Wake Vision test set. The Wake Vision (Combined) model is trained for longer than the VWW and Wake Vision (Quality) models but shows the value of using the training sets in unison even on the VWW test set.
4.3 Comprehensive Model Architecture Evaluation
To show that the benefits of using Wake Vision over VWW are not limited to MobileNetv2 architectures, we also train several other models on both training sets, Wake Vision (Quality) and VWW, and compare their performance. For this experiment, we use a collection of models commonly used in the TinyML community, such as models from the MCUNet [24] and Micronets [2] families. Furthermore, we include models from ColabNAS [11] to provide evaluation results on even smaller models.
Figure 3: Cross-evaluation results on the Wake Vision test set.
Each model is trained on each dataset using the Adam optimizer with a learning rate of 0.001 for 243,800 steps at a batch size of 512, equivalent to 100 epochs on the Wake Vision (Quality) training set. In addition, we apply random horizontal flips rotations and use 8-bit quantization-aware training. We use the checkpoint with the best validation score for testing.
In our comparison, we cross-evaluate each trained model’s performance on the test sets of both VWW and Wake Vision. Results on the Wake Vision test set can be found in Fig. 3, while the results for the VWW test set can be found in Fig. 5 alongside the complete data in Appendix A.
For the majority of models, the one trained on the Wake Vision (Quality) set achieves a higher accuracy on both the VWW and Wake Vision test set, which supports our claim that Wake Vision is a direct improvement over VWW. In some models, the VWW trained variant performs better on the VWW test set. This is only the case for the smallest models in our evaluation suite, which we believe may be due to these models being too small to generalize past the slight domain shift between the VWW and Wake Vision test sets. We additionally see that the accuracy discrepancy between identical models is often highest on Wake Vision’s Test set, further suggesting that the new test set is more challenging.
The boost in accuracy achieved by using the Wake Vision (Quality) Training set can be traded off for smaller model architectures (in some cases 3X fewer MACs) while preserving test accuracy. Accuracy and performance metrics, such as latency or model size, are in constant tension in TinyML models. Therefore, any boost in accuracy that does not impact the model’s architecture enables further efficiency.
4.4 Impact of Dataset Quality & Size in TinyML
Understanding the relationship between dataset quality and size is important for training efficient models in real-world scenarios. While gathering large amounts of data has become easier, ensuring high-label quality remains costly and time-consuming. This creates an important trade-off: is it better to have a smaller dataset with high-quality labels or a larger dataset with more label noise? The answer likely depends on the capacity of the model and the specific task.
We investigate the impact of dataset quality and size on Resnet [14] style models of varying capacity. We measure quality by the approximate rate of label errors. The Wake Vision (Quality) training set has an estimated error rate of around 7%. We simulate higher error rate datasets (15% and 30%) by flipping the binary label with a certain probability. Appendix E gives the derivation for the flip probability. We also sweep the dataset size by taking a slice of Wake Vision (Quality). Each model is trained for 50,000 steps on 224x224x3 images using AdamW [26], a learning rate of 0.001 with a cosine schedule and a weight decay of 0.004.
Figure 4 shows that smaller models (leftmost figure) are more sensitive to higher error rates than large models, similar to prior findings that model capacity can help memorize and ignore errors and outliers [40, 9, 13]. The lowest accuracy drop in the smallest model is 1. 3%, going from the base train (quality) error rate of around 7% to 15% compared to only a 0.5% accuracy loss for the largest model. In contrast, large models benefit more from big datasets and start to overfit on the smaller slices of the training data [18]. We also see the large slice of the 15% error rate dataset outperforming a smaller slice of the 7% error rate dataset.
Figure 4: Impact of dataset error rate and size on models of varying capacity. The $y$-axis is the Wake Vision test accuracy and the $x$-axis is the percentage of Wake Vision Train (Quality) used in training. The error % indicates the expected rate of incorrectly labeled samples.
4.5 Fine-Grained Benchmark Evaluation
Although aggregate metrics like overall accuracy or F1 score provide a high-level view of model performance, they mask critical failure modes and biases that emerge in real-world scenarios. For this reason, we included fine-grained benchmarks in Wake Vision (Section 3.3). Fine-grained evaluation across demographics, environmental conditions, and visual contexts helps us develop reliable TinyML Computer Vision systems for bespoke use cases. This is particularly important for person detection models deployed in real-world applications, where failures on certain subgroups or conditions could have serious implications for fairness and safety.
In Table 5, we compare the models from Sec. 4.2, trained on Wake Vision and VWW, across various scenarios to understand their relative strengths and limitations. The Wake Vision model exhibits superior robustness across the challenging settings exercised by our benchmarking suite, further demonstrating the effectiveness of the dataset. For instance, on the “Depictions” benchmark, which evaluates performance on images containing persons, non-person objects, or no depictions, the Wake Vision model achieves an F1 score of 0.71 for person depictions, outperforming the VWW model’s 0.66. Similarly, the Wake Vision model scores 0.86 for non-person depictions compared to the VWW model’s 0.82.
Table 5: Wake Vision Fine Grain Benchmark Suite. We report the samples in each set and the average F1 score across three Wake Vision models and three VWW models on the benchmark set.
Set Size F1-Score
Benchmark Val Test Wake Vision VWW
Gender Female $684$ $2157$ $0.93$ $0.89$
Male $1310$ $3918$ $0.91$ $0.88$
Unknown $1612$ $4940$ $0.77$ $0.78$
Age Young $275$ $884$ $0.94$ $0.90$
Middle $2133$ $6595$ $0.91$ $0.88$
Older $90$ $276$ $0.94$ $0.89$
Unknown $1299$ $3837$ $0.71$ $0.74$
Distance Near $5457$ 16 3331633316\,33316 333 $0.91$ $0.89$
Medium $2213$ $6876$ $0.85$ $0.84$
Far $398$ $1140$ $0.59$ $0.67$
Lighting Dark $3255$ $9420$ $0.85$ $0.81$
Normal 14 3151431514\,31514 315 43 0104301043\,01043 010 $0.82$ $0.82$
Bright $1012$ $3332$ $0.82$ $0.82$
Depictions Person $356$ $978$ $0.71$ $0.66$
Non-Person $352$ $1101$ $0.86$ $0.82$
No Depiction $8583$ 25 8022580225\,80225 802 $0.87$ $0.85$
The Wake Vision model also showcases improved performance on the “Age” benchmark, with F1 scores of 0.94, 0.91, and 0.94 for young, middle, and older individuals, surpassing the VWW model’s scores of 0.90, 0.88, and 0.89, respectively. This highlights the dataset’s effectiveness in enhancing model robustness for detecting people across various age groups, particularly the elderly demographic. In addition, the Wake Vision model demonstrates resilience to challenging lighting conditions, achieving F1 scores of 0.85 in dark lighting scenarios, compared to the VWW’s 0.81.
Notably, a VWW model only significantly outperforms a Wake Vision model on the far distance dataset at a score of 0.67 and 0.59, respectively. This is likely caused by a difference in how the two datasets are created. Where VWW filters out persons smaller than 0.5% of the image, this filter is set to 5% for Wake Vision. Therefore, the VWW dataset contains images with significantly smaller persons than the Wake Vision dataset, which can improve its performance on far away persons. We did experiment with setting the limit to 0.5% for Wake Vision however, we experienced that it disproportionately decreased the performance of models trained on the dataset in other settings.
Collectively, these results underscore the significance of Wake Vision’s large-scale, diverse data and fine-grained benchmarks, enabling the development of more robust and reliable TinyML person detection models across a wide range of realistic and real-world scenarios.
5 Ethical Considerations
The Wake Vision dataset and benchmark suite are designed to advance TinyML research by providing data resources and enabling detailed evaluation of person detection systems. While this work has significant potential benefits, we acknowledge several important ethical considerations:
Intended Benefits
Wake Vision can improve the quality of person detection systems that serve beneficial societal purposes, for example, an occupancy detection system that conserves energy by turning off the lights/HVAC/TV when a person leaves. The focus on TinyML and on-device inference helps protect user privacy by avoiding cloud-based processing of sensitive data.
Potential Misuse
We recognize that person detection technology can be misappropriated for harmful purposes, including weapons targeting or mass surveillance systems. While we cannot completely prevent misuse, we implement several safeguards, including releasing only evaluation sets, not training sets, for our fine-grained benchmark datasets.
Data Rights and Privacy
The images in Wake Vision are sourced from Flickr through Open Images [21, 19] under the CC-BY 2.0 license. While we have made extensive efforts to ensure proper licensing, we acknowledge the possibility that some images may have been uploaded without proper distribution rights. Furthermore, the dataset may contain potentially identifying biometric information, and given sufficient resources, individuals could potentially be identified. Due to the scale of the dataset, exhaustive manual verification for offensive content or privacy concerns is infeasible.
Fairness and Representation
Our benchmark suite is specifically designed to evaluate model fairness and robustness across different demographics and conditions. However, we acknowledge that these same demographic benchmarks could be misused to develop systems that classify gender and age without consent, potentially enabling privacy violations or discriminatory practices. We have carefully balanced these concerns against the importance of ensuring that person detection systems work equitably for all populations.
6 Conclusion
The lack of large, high-quality datasets has been a significant bottleneck in TinyML research. Through Wake Vision, we address this challenge by providing a comprehensive dataset and benchmark suite that advances the field in several key dimensions. Wake Vision surpasses the current state-of-the-art VWW dataset by offering  100x more images while maintaining ease of use and improving label accuracy. Our experiments show that Wake Vision serves as an effective drop-in replacement for VWW, consistently improving performance across various model architectures and training scenarios. Moreover, our investigation into dataset quality versus quantity trade-offs reveals important insights specific to TinyML systems. Unlike large models that can effectively learn from noisy data, we find that TinyML models are particularly sensitive to label quality because of their limited capacity. However, we also demonstrate that carefully designed training strategies—such as our approach of pre-training on Wake Vision (Large) followed by fine-tuning on Wake Vision (Quality)—can effectively leverage both quantity and quality to achieve superior performance. Finally, the Wake Vision benchmark suite, with its fine-grained evaluation sets, establishes new standards to evaluate the robustness of the TinyML model in demographics, environmental conditions, and visual contexts. These benchmarks enable researchers to identify and address potential biases and failure modes that could otherwise go undetected using traditional evaluation metrics.
Appendix A Additional Benchmark Results
Table Tab. 6 provides input shape, Flash, RAM, and MACs for each model presented in section 4.2 as well as the mean and standard deviation of test accuracies for both Wake Vision and VWW datasets. RAM and Flash have been measured using the “stm32tflm” utility of X-CUBE-AI 8.1.0, whereas MACs with “stm32ai”. Links pointing to the original models used for training are also present for each model family involved in the experiment.
Figure Fig. 5 shows cross-evaluation results on the Visual Wake Words test set. The majority of models trained on the Wake Vision dataset obtain better test accuracy on the VWW than the ones trained on the VWW. Only the smallest models do not achieve a higher test accuracy when trained on the Wake Vision, likely due to the lack of enough resources to overcome the slight domain shift between the two datasets.
Model input RAM Flash MACs Test Set Training Set
.tflite volume [kiB] [kiB] [MM] WV (Qual.) VWW
MobileNetV2  [^1] https://keras.io/api/applications/mobilenet/[33]
MobileNetV2_0.25 (224,224,3) 1,244.5 410.55 36,453,732 WV $84.9\pm 0.11$ $83.8\pm 0.23$
VWW $88.6\pm 0.17$ $88.3\pm 0.29$
MCUNet  [^2] https://github.com/mit-han-lab/mcunet/blob/master/mcunet/model_zoo.py[24]
10fps_vww (64,64,3) 168.5 533.84 5,998,334 WV $81.7\pm 0.28$ $77.6\pm 0.31$
VWW $81\pm 0.12$ $80.8\pm 0.1$
5fps_vww (80,80,3) 226.5 624.55 11,645,502 WV $82.9\pm 0.29$ $79.6\pm 0.26$
VWW $83.2\pm 0.57$ $82.6\pm 0.16$
320kb-1mb_vww (144,144,3) 393 923.76 56,022,934 WV $85.6\pm 0.34$ $82.4\pm 0.62$
VWW $86.5\pm 1.01$ $86\pm 0.25$
Micronets [^3] https://github.com/ARM-software/ML-zoo/tree/master/models/visual_wake_words[2]
vww2_50_50_INT8 (50,50,1) 71.50 225.54 3,167,382 WV $71.9\pm 0.67$ $66.2\pm 0.88$
VWW $70.8\pm 0.68$ $70.6\pm 0.64$
vww3_128_128_INT8 (128,128,1) 137.50 463.73 22,690,291 WV $77.8\pm 0.56$ $72.6\pm 1.01$
VWW $78.3\pm 0.91$ $77.6\pm 1.13$
vww4_128_128_INT8 (128,128,1) 123.50 417.03 18,963,302 WV $77.9\pm 0.6$ $71.3\pm 1.03$
VWW $78.4\pm 0.51$ $76.5\pm 0.37$
ColabNAS444 https://github.com/AndreaMattiaGaravagno/ColabNAS/tree/main/models [11]
k_2_c_3 (50,50,3) 18.5 7.66 250,256 WV $70.6\pm 0.96$ $69.3\pm 0.97$
VWW $65.6\pm 0.66$ $70.7\pm 0.08$
k_4_c_5 (50,50,3) 22 18.49 688,790 WV $75.7\pm 0.18$ $74\pm 0.23$
VWW $69.9\pm 0.26$ $75.5\pm 0,64$
k_8_c_5 (50,50,3) 32.5 44.56 2,135,476 WV $77.3\pm 0.37$ $75\pm 0.15$
VWW $73\pm 0.91$ $77.3\pm 0.57$
Table 6: Raw data of reconstructed models.
Figure 5: Cross-evaluation results on the Visual Wake Words test set
Appendix B Code Repository
The code used to generate the Wake Vision dataset is available at the following GitHub repository: https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
This repo contains the code to generate Wake Vision and the benchmark suite, as well as the code to train and evaluate models. This code is sufficient to reproduce all results in the paper.
Appendix C Flowchart of Bounding Box Filtering Process
Fig. 6 illustrates the filtering process for Wake Vision when using Open Images’ bounding box labels as the label source.
Figure 6: A flowchart of the filtering of an image according to default Bounding Box filtering rules.
Figure 7: Effects of scaling the image size vs. the model width on Wake Vision test accuracy and the Distance-Far benchmark set. The Distance-Far F1 score (right) is far more sensitive to changes in the image size than the overall test accuracy (left). Image sizes refer to square image resolutions (e.g. 96x96) and model size refers to the MobileNetV2 width multiplier. When not specified the image size is 160 and the width multiplier is 1.0.
Appendix D Model Design Case Study
Basic test set performance can often misrepresent a model’s performance, given the typical domain shift between images scraped from the internet and real-world use cases. This issue is exacerbated when ML practitioners must trade off accuracy for model performance and size, which is necessary for TinyML use cases. A design decision might have seemingly little impact on the test accuracy but may destroy real-world performance depending on the deployment environment. For example, a person detection system may only operate in dark lighting conditions, but the test dataset has an insignificant number of dark samples; therefore, the test accuracy will not reflect the real-world accuracy.
The benchmark suite enables more holistic analysis during the design phase. To show this use we perform a series of scaling experiments employing typical TinyML compression techniques and identifying under which circumstances these techniques are appropriate. While these results can inform ML practitioners, our intention is to demonstrate the usefulness of the benchmark suite.
D.1 Image Size vs. Model Width Scaling
We train two series of MobileNetV2 models: one series that sweeps the input image size [64-256] and one that sweeps the width multiplier of a model [0.1-1.5]. We then benchmark these models on the Wake Vision test set as well as the far distance benchmark. We plot these results against the number of multiply-accumulate (MAC) operations in the model as a proxy for on-device latency [2].
The results in Table 7 (left) show that when looking exclusively at the high-level metric (i.e., test accuracy), scaling the input image size has a similar impact as scaling the model size. However, as shown in Figure 7 (right), when we consider only samples where the person is far away from the camera (i.e., the distance benchmark), we observe a much more significant impact when scaling the image size. In the case of distant subjects, the image size becomes the bottleneck.
These findings suggest that for ML developers targeting use cases where the subject is likely to be far from the camera, prioritizing larger input image sizes over wider models may be more beneficial. However, this critical design consideration could be obscured when solely relying on high-level metrics like overall test accuracy. The distance benchmark in Wake Vision effectively unveils the disproportionate impact of image size on model performance for distant subjects, enabling more informed decision-making during model optimization.
D.2 Quantization
Quantization is a crucial technique for deploying efficient TinyML models, offering substantial benefits in terms of reduced latency, memory footprint, and model size. However, prior work has suggested that quantization can disproportionately impact the performance of models on underrepresented subsets of data [17]. To assess the implications of quantization in the context of person detection, we investigate the impacts of int8 quantization on a model’s benchmark results across Wake Vision’s fine-grained benchmarks.
Our findings show negligible degradation in performance across all benchmarks (±$\pm 0.004$ F1) when employing int8 quantization, even on outlier sets. This result contradicts the previously observed disproportionate impact of quantization on underrepresented subsets. We speculate that person detection may be a relatively simple task, potentially explaining why we do not observe this specific property of quantization in our experiments. Given the negligible performance degradation and the substantial latency, memory, and model size benefits of quantization, we conclude that quantization is a win for person detection.
D.3 Grayscale
Converting a model’s input image channels to grayscale from RGB is a commonly employed optimization in the TinyML field [2] as it can substantially reduce a model’s memory consumption. We observed, however, that the grayscale optimization disproportionately impacts images on the brighter end of the spectrum as illustrated in Figure 8. This further demonstrates the importance of fine-grained analysis, as some real-world deployment environments might be far brighter than the average Wake Vision test sample.
Figure 8: Impact of grayscale input images on the lighting benchmarks: dark, normal light, and bright. Models that use grayscale input images are more sensitive to bright lighting conditions than RGB.
Appendix E Inducing Label Errors
The goal is to make a single pass through the dataset and flip the labels of a binary classification dataset such that the expected label error rate is $d$. There is an underlying rate of label errors $e$. If we flip one of these underlying errors, we correct it, thereby inadvertently decreasing the label error rate. After flipping labels with a probability of $p$ we can claim that the likelihood of a single label being correct is the probability that we flipped the label and it was originally an error plus the probability that we didn’t flip the label and it was originally correct: $1-d=p*e+(1-p)(1-e)$. Then solving for $p$ gives $p=(e-d)/(2e-1)$. A current flaw of this method is that the injected label errors are not consistent between epochs, which would likely be less destructive to a model’s accuracy since the same errors are not reinforced each epoch. This could also potentially explain why large models in Fig. 4 don’t overfit on training data with higher label errors, as the inconsistent label noise has a regularizing effect.
Appendix F Fine Grained Benchmark Images
Fig. 9 gives example images for each of the benchmarks in the suite.
Gender
Female
Male
Gender Unknown
Age
Young
Middle
Older
Age Unknown
Distance
Near
Medium
Far
Lighting
Dark
Normal
Bright
Depictions
Person
Non-Person
No Depiction
Figure 9: Images from each fine grained benchmark dataset.
Appendix G Open Image Label Distribution
Wake Vision is derived from Open Images V7, therefore the diversity of subjects in the images should follow the distribution of labels in Open Images. The Label distribution of Open images can be found here: https://storage.googleapis.com/openimages/web/factsfigures_v7.html#statistics
Appendix H Dataset Access and Organization
Wake Vision is available online via TFDS and Hugging Face Datasets. We exclude direct links to the dataset during the review period to avoid breaking anonymity.
The dataset is organized as a set of compressed tar files containing images and a series of label CSVs. The label CSVs are organized such that the file name of the image is the identifying index. CSVs for all dataset splits include the person and depiction label. The Validation and Test label CSVs also have flags that denote a sample’s inclusion into a fine-grain benchmark set (e.g. Distance-Near). This structure makes it easy to access just the required data without requiring a full download. It also ensures the dataset can be easily updated as new versions are introduced.
Appendix I Label Generation Details
Person Labels. The most straightforward Open Images label classes to label as a person in Wake Vision are the "person" label and its subcategories (Configurable, but defaults listed in Appendix L). All of these are relabelled as persons in Wake Vision. These labels are present as both image-level labels and bounding box labels.
We furthermore inspect the image-level label classes for synonyms and umbrella terms for all the person related labels. This search resulted in an additional six person related label classes. These person related label classes will only be used in the image level label configuration.
Lastly, the image level person labels in Open Images have an associated confidence score. This score ranges from 0-10. Labels that have been verified by humans to be absent from an image have a score of 0, while labels that have been verified present by humans have a score of 10. Purely machine-generated labels have a fractional confidence score that is generally $>=5$ [21, 19]. By default, we only respect labels that have a minimum confidence of 7. Labels below this threshold are ignored.
Person Body Part Labels. Body parts are more challenging to relabel, as it is dependent on the use case whether a body part should be considered a person.
For example a camera that detects whether a person is inside a room to decide if the light should be switched on would want to consider body parts as a person, as this will keep the lights on even when the person is only partly in the camera frame. For waking up electronics, however, it may not make sense to consider body parts as persons. This could, e.g., mean that a computer would turn on when detecting a foot.
To cater to both use cases we include a flag in our open-source dataset creation code to set whether body parts should be considered persons. By default we consider body parts as persons.
Depictions. Open Images bounding box labels contain metadata about whether an object is a depiction, e.g., a painting or a photograph of a person. This presents the challenge of how to handle depictions. While most use cases would not consider a depiction a person, it could make sense to either exclude them to make training easier, or include them as non-persons to make a model resistant to seeing depictions when deployed.
In line with how we handle body parts, we therefore include a flag for our open-source dataset creation code to set whether depictions should be excluded or considered non-persons. By default depictions are considered non-persons.
Bounding Box Size. The VWW dataset only considered a Common Objects in Context (COCO) person to be a person if the bounding box around the person took up at least 0.5% of the image [5]. If the person took up less than 0.5% of the image, the image was excluded from the dataset. For Wake Vision, we observed a decrease in performance at the 0.5% limit. We theorize that the cause of this decrease is that a person becomes indistinguishable at such a low percentage of already small images. Therefore we adopted a 5% threshold for the Wake Vision dataset instead. For different requirements, users can change a configuration parameter in our open-source dataset creation code.
Appendix J Standardized Evaluation
VWW has no standardized way to evaluate performance on the dataset. This makes it challenging to compare works based on the dataset, since performance difference can come down to choices outside the contribution of the work. e.g., two works could be contributing with model improvements, but use different pre-processing pipelines that skews results.
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec. 3.3.
Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier [^5] Also known as alpha. of 0.25 [33]. For model-centric contributions we propose the following pre-processing pipeline:
  1. 1.
    Cast image pixel value datatype from 8-bit integers into 32-bit floating points
  2. 2.
    Resize image such that the shortest side matches the model input size
  3. 3.
    Perform a center crop on the image such that the longest size matches the model input size
  4. 4.
    Normalize pixel values to between -1 and 1 sample wise
  5. 5.
    Use image tensor as input features and person label as target feature
Appendix K Open Images Download
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab. 7.
Table 7: Number of images downloaded from Open Images v7. Download occurred between the 28th of November to the 5th of December
Train Validation Test
Downloaded 7,936,979 36,406 109,305
Errors 1,055,669 5,214 16,131
Appendix L Person Label Classes
In our default configuration, we consider the following Open Images v7 labels to be a person for the Wake Vision dataset:
  • Person
  • Woman (Subcategory of Person)
  • Man (Subcategory of Person)
  • Girl (Subcategory of Person)
  • Boy (Subcategory of Person)
  • Human body (Part of Person)
  • Human face (Part of Person)
  • Human head (Part of Person)
  • Human (Person synonym - Only in Image Level Label Configuration)
  • Female person (Woman synonym - Only in Image Level Label Configuration)
  • Male person (Man synonym - Only in Image Level Label Configuration)
  • Child (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
  • Adolescent (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
  • Youth (Umbrella term for Girl & Boy - Only in Image Level Label Configuration)
The following labels are considered body parts and are labelled according to the body part configuration.:
  • Human eye (Part of Person)
  • Skull (Part of Person)
  • Human mouth (Part of Person)
  • Human ear (Part of Person)
  • Human nose (Part of Person)
  • Human hair (Part of Person)
  • Human hand (Part of Person)
  • Human foot (Part of Person)
  • Human arm (Part of Person)
  • Human leg (Part of Person)
  • Beard (Part of Person)
Appendix M Wake Vision Dataset Size
See Tab. 8.
Table 8: Amount of images in the Wake Vision dataset
Person Images Non-Person Images Excluded
Wake Vision (Large) Training Dataset 2,880,214 2,880,214 2,176,551
Wake Vision (Quality) Training Dataset 624,115 624,115 6,688,749
Validation Dataset 9,291 9,291 17,824
Test Dataset 27,881 27,881 53,543
Appendix N Manual Label Correction
We use the Scale Rapid tool from Scale AI to relabel the Wake Vision validation and test sets. Figure 10 shows a screenshot of the labeling menu presented to the human labelers. The labelers were instructed to label an image as a "person" if a person was present anywhere in the image and "no person" if no visible person was present. The labelers also indicated if the image contained a depiction of a person or a "picture of a picture/reflection", which we use for metadata.
The cost per image was \$0.10 USD and is set based on Scale’s pricing structure. Each image is labeled by 3 different labelers to form a consensus. The total cost of labeling was \$7089.8 USD. The authors do not know the hourly rate paid to each labeler, but Scale lists \$18/hr on job postings at the time of writing. For context on the difficulty of the task, the authors were able to average around 500 images per hour at a reasonable pace when estimating error rates.
Figure 10: A screenshot of the labeling menu used to manually label the validation and test sets.
Appendix O Automatic Label Correction
To address the challenge of correcting Wake Vision validation and test labels we initially employed the Confident Learning technique [29] to intelligently identify potential label errors. We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [30]. Confident Learning flagged suspected mislabeled instances, which we inspected and corrected through a manual verification process.
Table 9: Amount of label errors identified and corrected using Confident Learning.
Dataset Total Suggested Corrected Suggestion
Split Size Errors Errors Accuracy
Validation 18,582 632 81 12.82%
Test 51,282 1672 267 15.97%
As shown in Table 9, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets. However, only between 12 and 16% of these possible label errors were legitimate errors. We corrected these identified label errors in the Wake Vision validation and test sets.
Given this low acceptance rate of label issues identified by Confident Learning, we concluded that we could not automate label cleaning to a point where no human in the loop is needed. This made this strategy too human-intensive to be applied to the much larger training sets.
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in Sec. 3.2 and deprecate the automatic label corrections described in this section.
Appendix P Author Statement
The authors of this paper hereby confirm that we bear all responsibility in case of violation of rights, etc, and confirm the data license to be CC BY 4.0 for our labels, and that all images in the dataset have been uploaded to Flickr with a CC BY 2.0 license. Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
Appendix Q Hosting, License and Maintenance Plan
The dataset is hosted on HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images using our open-source filtering code. The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
Wake Vision’s labels and metadata are licensed under a CC BY 4.0 license. Images in the dataset have all been uploaded to Flickr with a CC BY 2.0 license. Note: while the Open Images authors tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image and users should verify the license for each image themselves.
Appendix R DOI and Croissant URLs
The DOI and URLs will be provided after the review period to avoid breaking anonymity.
Appendix S Datasheet for Wake Vision
This document is based on Datasheets for Datasets by Gebru et al. [12]. Please see the most updated version here.
{mdframed}
[linecolor=black]
MOTIVATION
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
Wake Vision was created as a prototype for a new generation of datasets for TinyML. Where previous TinyML datasets were small due to limited resources, Wake Vision is a production grade dataset consisting of almost 6 million images and with an extensively cleaned validation and test set. Wake Vision specifically targets the binary person classification task. Such a large dataset enables TinyML researchers to explore questions previously impossible. For example, the size of the Wake Vision dataset enables new research into the tradeoff that often exist between dataset size and sample quality.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
These details will be provided after the review period to avoid breaking anonymity.
What support was needed to make this dataset? (e.g.who funded the creation of the dataset? If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
These details will be provided after the review period to avoid breaking anonymity.
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
COMPOSITION
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
The dataset is composed of images, labels, and metadata. The images are sourced from Open Images. The labels are binary, with ’1’ indicating there is a person somewhere in the image and ’0’ indicating that there is not a person in the image. The metadata is used to form the benchmark suite where we can evaluate models in specific challenging settings. Each image has a binary variable for each fine-grained benchmark where a ’1’ indicates that the image belongs to the fine-grained benchmark set.
How many instances are there in total (of each type, if appropriate)?
There are over 6 million labeled images in the dataset.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The dataset uses a large subset of the Open Images dataset (6 million out of a possible 9 million images). We use a subset in order to have a balanced number of person samples and non-person samples. We also filter the dataset for quality. The filtering removes subjects that are too far and outside of a center crop. Open Images sources images from Flickr that are licensed as CC-BY 2.0. This is likely not a fully representative sample of the overall world population and likely skews heavily towards a European or North American population. See the Open Images paper [21] for more information.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
The dataset is composed of images, labels, and metadata.
Is there a label or target associated with each instance? If so, please provide a description.
Yes, every image has a binary label to indicate if a person is visible in the image. Images in the some dataset splits include more metadata labels, e.g., about distance of subject or whether the subject is a depiction.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
The training set does not contain the metadata used for the fine grained benchmark suite. The dataset contains two training sets, one of higher quality and one larger. The higher quality dataset and other sets contain information about depictions and body parts. The large training set does not contain this information.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
No
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
Yes there are separate splits for training, validation, and testing. These splits are generated from equivalent splits in Open Images. The dataset contains two training splits, one focusing on quality of samples and one on number of samples.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
Yes there are label errors in the dataset due to the labeling process. We discuss this at length in the paper.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
The dataset is self-contained. We re-host the images, labels, and metadata. The images may be subject to different licenses than our labels and metadata. All images have originally been hosted on Flickr under a permissive CC BY license.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)? If so, please provide a description.
No, the images were posted online with the CC-BY 2.0 license. It is possible an image was posted without the individual’s consent. While we tried to identify images that are licensed under a Creative Commons Attribution license, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No, not to our knowledge. Due to the size of the dataset we have not had the resources to manually verify this.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Yes, the dataset contains images of people.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
Yes the benchmark suite contains subsets that focus on perceived age and gender. These labels have been sourced from the More Inclusive Annotations of People [34]. The purpose of these benchmarks are to test the fairness of a trained model across these subpopulation and we do not condone using these labels to train a model. The size of these sets is listed in Table 5.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
It could be possible to identify a person by their face or other visible characteristics in the image. We do not share any other personal information besides the images.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
No, not to our knowledge. Some personal or sensitive information may be visible in the images. However, the images were shared publicly under a CC-BY 2.0 license. We do not include any personal information beyond the images.
Any other comments?
The images used were published publicly under a CC-BY 2.0 license and are already used in the Open Images [19] dataset and its derivatives. The authors of Open images tried to identify images that are licensed under a Creative Commons Attribution license but make no representations or warranties regarding the license status of each image and a user should verify the license for each image themselves.
{mdframed}
[linecolor=black]
COLLECTION
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The dataset is derived from Open Images [19], which in turn acquired its images from Flickr. Section 3.1 describes the process.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. Finally, list when the dataset was first published.
The original data collection was conducted in Noember of 2015. The Open Images paper [21] details the image acquisition process.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
The images were sources from Flickr in an automated pipeline. The original labeling pipeline was a combination of automated labeling and manual verification and labeling. The Open Images paper details this process [21]. We detail our process for relabeling and filtering the dataset in Section 3.1.
What was the resource cost of collecting the data? (e.g. what were the required computational resources, and the associated financial costs, and energy consumption - estimate the carbon footprint. See Strubell et al.[35] for approaches in this area.)
We used cloud TPU credits provided by the Google Cloud TRC program. The primary source of required resources was model training to evaluate the dataset and the storage and bandwidth required to process and upload the dataset to hosting locations. The total size of the dataset is approximately 2 TB. To the authors knowledge, there is no public information about the power draw of TPUs. This matches the conclusions drawn by Strubell et al.[35]. This makes it impossible to disclose the energy consumption and estimated carbon footprint of the project.
We used paid crowdsource labelers to relabel the validation and test sets. The total cost of labeling was approximately \$7,000.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
The dataset is a large subset of the total Open Images dataset. The subset is generated by a deterministic filtering process described in Section 3.1.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
We used paid crowdsource labelers to relabel the validation and test sets. The total cost of labeling was approximately \$7,000. We used the Scale Rapid labeling service. A total of 70,898 images were labeled at a cost of \$0.1 per image and \$0.033 per individual label. Appendix 3.2 gives additional detail.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No
Does the dataset relate to people? If not, you may skip the remainder of the questions in this section.
Yes the dataset contains images of people.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
We obtained the images via the Open Images dataset [21] which originally sourced its images from Flickr [10].
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
No the individuals were not notified to our knowledge. The images were posted publicly under the CC-BY 2.0 license.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes, the images were posted publicly under the CC-BY 2.0 license. It is possible that the images were uploaded by a third party without the consent of the individual in the image.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate)
No we do not have a mechanism for an individual the revoke their consent. As with Open Images, we make no representations or warranties regarding the license status of each image, and a user should verify the license for each image themselves via the Flickr link in the metadata. Note that it is impossible to revoke a CC BY license after publication [23].
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis)been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
No formal analysis has been conducted. We describe the ethical considerations of the dataset in Section 5.
Any other comments?
No additional comments.
{mdframed}
[linecolor=black]
PREPROCESSING / CLEANING / LABELING
Was any preprocessing/cleaning/labeling of the data done(e.g.,discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
Yes we filtered the data as shown in Figure 6.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
The "raw" data is accessible in the Open Images dataset.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Yes, we open-source our filtering code. https://anonymous.4open.science/r/Wake_Vision-DD19/README.md
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
USES
Has the dataset been used for any tasks already? If so, please provide a description.
The dataset has been used for the experiments reported in the paper.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
The usage of the dataset can be tracked via the citations of this paper. We plan to host competitions using the dataset and will track participants.
What (other) tasks could the dataset be used for?
The dataset can be used as a benchmark for TinyML models and as the basis for data-centric AI research given the prevalence of label errors in the training set and the cleaned nature of the validation and test sets.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
Yes, the estimated error rates are reported in Table 2. We also discuss the high likelihood of false negatives in Wake Vision’s training sets in Section 3.2. These errors are due to the labeling and filtering process. We mitigate the impact of these errors by manually labeling the validation and test sets as well as study their impact for TinyML models in Section 4.4.
Are there tasks for which the dataset should not be used? If so, please provide a description.
The perceived age and gender benchmarks should not be used to train a gender or age classifier.
Any other comments?
No additional comments
{mdframed}
[linecolor=black]
DISTRIBUTION
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
The dataset is hosted on Harvard Dataverse and HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code. We plan for the TinyML foundation (https://www.tinyml.org/) to assume responsibility for hosting and maintaining the dataset long-term.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
Yes, the dataset is hosted on Harvard Dataverse and HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open source filtering code.
When will the dataset be distributed?
The Dataset is publicly accessible at the time of writing.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The dataset is published under the CC-BY 4.0 license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No
Any other comments?
No additional comments.
{mdframed}
[linecolor=black]
MAINTENANCE
Who is supporting/hosting/maintaining the dataset?
The dataset is hosted on HuggingFace Datasets. Additionally, the dataset is available through TensorFlow datasets and can be regenerated from Open Images with our open-source filtering code. The Edge AI Foundation will assume responsibility for hosting and maintaining the dataset long-term.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Removed during review to maintain anonymity.
Is there an erratum? If so, please provide a link or other access point.
No
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
The dataset may be updated if any problematic instances are discovered or the label errors are corrected. In this case, we will issue a new version of the dataset to ensure past results are contextualized under the version of the dataset that was used.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
There is no limit.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Yes, we will continue to host older versions of the dataset. Details will be provided after review to preserve anonymity.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
Yes, the dataset is licensed permissively, and the code used to create Wake Vision is open source.
Any other comments?
No additional comments